The head and tail of train moving with a constant acceleration cross an electric pole on the track  side with velocities u and v respectively. Find the velocity with which the middle point the train crosses the pole.

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
 

v^2 - u^2 = 2as

Find w such that

w^2 - u^2 = 2 (a)(s/2)

 

Eliminate (as) between the two equations.

 

w = sqrt((v² + u²)/2)